OpenStudy (landon34):

Which of the following expressions represents the least common multiple of 15, 18, and 25? 2*3*5 2*3^2*5^2 2 · 3^3 · 5^3 2 · 3^2 · 5^3

OpenStudy (whovianchick):

Well I figured if you're asking a question you need help XD gimme a sec

OpenStudy (welshfella):

15 = 3 * 5 18 = 2*3*3 25 = 5*5 where there are dupicates between the factors use the number once otherwise use all numbers

OpenStudy (welshfella):